# Weakly-Supervised 3D Visual Grounding based on Visual Linguistic Alignment

### v3

@@ IV-C2 Investigating the Influence of Top-$k$ Query Category Predictions in Proposal Filtering
We investigate the influence of using different top-$k$ query category predictions in our category-oriented proposal filtering strategy. The experiments are conducted on the Nr3D dataset, and the results are shown in Table. V. We set $k$ in four different values, i.e., $k\in\{1,2,3,4\}$. It can be observed that keeping more query category predictions brings higher accuracy, which shows that keeping more possible categories from the query could provide more semantic information to filter invalid 3D proposals, and is helpful to 3D visual grounding. We take $k=3$ as the default setting since a value of $k$ that is too large might introduce an excess of interference candidates, leading to a negative impact on the network’s performance.
IV-C3 Further Analysis
@@
We dicuss the extension of the proposed method and design some other baselines to further analysis the model: (a) Random: randomly selects a proposal as the target proposal of text query. (b) Hard Weakly Supervised: uses CLIP to compare 2D image regions and text queries, and leverages their matching results as the pseudo-labels for 3D proposals and text queries. (c) Fully Supervised: removes the 2D branch of 3D-VLA, and directly utilizes 3D labels for fully supervised training. The results are provided in Table.VI. We find that our 3D-VLA is better than the Hard Weakly Supervised since pseudo-labels might not be accurate and disturb the model, and our method is more robust via leverages natural 3D-2D correspondence for efficiency embedding learning. For Fully Supervised, although ours doesn’t completely outperform it across all the subsets, the performance is still comparable.
IV-C4 Qualitative Results
@@ IV-D Limitation
There are several limitations of our work and still much to do to realize the full potential of the proposed approach. We still follow Wang et al[63] and employ the pretrain model to extract the proposals in advance. Therefore, as shown in Fig. 6, the performance of our method is largely limited by the accuracy of the pretrained detection model. Besides, our method still require extra 2D image during training so that it can not be applied for those datasets only with 3D point cloud. Using rendering image technology to generate high-quality 2D synthetic images may be a good solution to deal with this problem. These limitation are direct avenues for future work.
@@ V Conclusion
In this paper, we propose to tackle the weakly supervised 3D visual grounding from a novel perspective towards Visual Linguistic Alignment, in an effort to address the shortage of object-sentence annotations. Specifically, our 3D-VLA leverages the superior ability of current advanced VLMs to align the semantics among texts and 2D images, as well as the naturally existing correspondences between 2D images and 3D point clouds, such that implicitly constructing correspondences between texts and 3D point clouds. During 3D-VLA inference, we exploit the learned text-3D correspondence to help ground the text queries to the referred 3D objects without regarding to 2D images. Through the designed scheme, a significant breakthrough is achieved than previous works, and the advantage of our 3D-VLA are also analyzed in detail. We believe these analyses can provide valuable insights to facilitate the future research of weakly supervised 3D visual grounding.

### v4

@@ IV-B1 ScanRefer
We also compare our 3D-VLA with zero-shot 3D visual grounding methods. For OpenScene [67] and LERF [66], we follow the methodology in LLM-Grounder [68] and apply DBSCAN clustering [69] to points with high cosine similarity between the point cloud and text embeddings. We then draw bounding boxes around these clustered points to identify target objects. Despite the zero-shot capabilities of these large models, our 3D-VLA consistently outperforms them across all subsets on ScanRefer, as shown in Table. I. This is primarily due to our model’s ability to leverage category information from each query, which provides prior knowledge to improve object localization. Unlike zero-shot methods, which generally lack specialization in indoor environments, our task-aware classification architecture allows the model to transfer indoor-specific embeddings, enhancing its understanding of such scenes. Additionally, our category-oriented proposal filtering isolates relevant objects by minimizing interference from irrelevant ones, further boosting the model’s localization accuracy.
@@ IV-C2 Investigating the Influence of Top-$k$ Query Category Predictions in Proposal Filtering
We investigate the influence of using different top-$k$ query category predictions in our category-oriented proposal filtering strategy. The experiments are conducted on the Nr3D dataset, and the results are shown in Table. VI. We set $k$ in four different values, i.e., $k\in\{1,2,3,4\}$. It can be observed that keeping more query category predictions brings higher accuracy, which shows that keeping more possible categories from the query could provide more semantic information to filter invalid 3D proposals, and is helpful to 3D visual grounding. We take $k=3$ as the default setting since a value of $k$ that is too large might introduce an excess of interference candidates, leading to a negative impact on the network’s performance.
IV-C3 Investigating the influence of potential inaccuracies in the 2D-3D correspondences on overall performance
It is well known that datasets may have potential inaccuracies in 2D-3D correspondences which is more likely to occur when the objects are small or the scene is complex. In such cases, the projected object may be too small or its location in the 2D image may be difficult to determine, leading to boundary misalignments and potential cutting errors. Additionally, misalignment of points along the object’s boundary can result in an overly small 2D area, further disrupting the localization.
Therefore, We conduct experiments using both the original, unmodified 2D projection regions and our boundary-extended approach to investigate the influence of potential inaccuracies in the 2D-3D correspondences on overall performance. The term ”Unmodified Projection” refers to the method using unmodified projected areas, while the term ”Boundary-Extended Projection” refers to the method using the expanded projection, as mentioned in Sec.III-A, where the projected 2D bounding box $[x,y,w,h]$ is expanded by 10% to produce the final partition area $[x,y,w+0.2*w,h+0.2*h]$ to avoid the potential inaccuracies in the 2D-3D correspondences. The results in Table VIII show that the performance of our boundary-extended approach is comparable to that of the original 2D projection regions, indicating that the dataset’s projection accuracy is generally reliable.
IV-C4 Investigating the computational cost
(a) Data Preparation Time: Prior to training, our method requires projecting the 3D point cloud to 2D to identify corresponding regions. While this projection is computationally intensive, we mitigate this by pre-computing the projections offline, significantly reducing the time burden during training (see Tab. VIII for details). (b) Computational Complexity: Our model is trained on a V100 GPU, leveraging PointNet++ as the underlying 3D architecture, which contributes to its lightweight nature. Our model requires approximately 14GB of GPU memory with batch size 48 for training and 3GB of GPU memory with batch size 48 for inference in Referit3D dataset, demonstrating its efficiency compared to other more resource-intensive methods. (c) Inference Time: In Tab. VIII, we present the inference runtime of our method. While our model does not yet achieve real-time speeds, its inference time remains competitive, allowing for practical deployment in applications where real-time performance is not critical. We are also actively exploring optimizations to further enhance inference speed.
IV-C5 Further Analysis
@@
We further analyze the performance of our model by designing several additional model variants: (a) Random Proposal Selection (RPS.): randomly selects a proposal as the target proposal for the text query; (b) CLIP-Based Weak Supervision (CBWS.): uses CLIP to compare 2D image regions and text queries, leveraging their matching results as pseudo-labels for 3D proposals and text queries; (c) Ground-Truth Supervision (GTS.): removes the 2D branch of 3D-VLA, and directly utilizes 3D labels for fully supervised training. The results are provided in Table VI. We find that our 3D-VLA method outperforms the CLIP-Based Weak Supervision method, as pseudo-labels may be inaccurate and hinder the model’s performance. Moreover, our method is more robust, leveraging natural 3D-2D correspondences for efficient embedding learning. For the Ground-Truth Supervision baseline, although our method does not always outperform it across all subsets, the performance remains comparable, which demonstrates the efficacy of our approach.
IV-C6 Qualitative Results
@@
IV-C7 Can 3D-VLA generalize to outdoor scenes with diverse lighting and object scales?
While our method has been primarily evaluated on indoor datasets, we believe it has strong potential for generalization to outdoor environments. The task-aware classification architecture of our 3D-VLA is designed to adapt to diverse environmental conditions, suggesting that it should, in theory, be capable of effectively localizing target objects in outdoor scenes, irrespective of lighting variations or object scale.
@@ IV-D Limitation
There are several limitations of our work and still much to do to realize the full potential of the proposed approach. Firstly, We still follow Wang et al[63] and employ the pretrain model to extract the proposals in advance. Therefore, as shown in Fig. 6, the performance of our method is largely limited by the accuracy of the pretrained detection model. Secondly, our method still require extra 2D image during training so that it can not be applied for those datasets only with 3D point cloud. Using rendering image technology to generate high-quality 2D synthetic images may be a good solution to deal with this problem. Besides, when multiple similar objects are placed next to each other and the query involves a relation like ”next,” the model may struggle to disambiguate between the objects. This issue is not unique to our 3D-VLA; even fully supervised methods face challenges with ambiguous relational queries. Lastly, We recognize the potential benefits of integrating zero-shot learning techniques, especially those using large language models (LLMs) like GPT-4. Models such as LLM-Grounder leverage environmental context and relational information for better object localization. We believe incorporating these techniques into our weakly supervised framework could enhance performance. These limitation are direct avenues for future work.
@@ V Conclusion
In this paper, we propose to tackle the weakly supervised 3D visual grounding from a novel perspective towards Visual Language Alignment, in an effort to address the shortage of object-sentence annotations. Specifically, our 3D-VLA leverages the superior ability of current advanced VLMs to align the semantics among texts and 2D images, as well as the naturally existing correspondences between 2D images and 3D point clouds, such that implicitly constructing correspondences between texts and 3D point clouds. During 3D-VLA inference, we exploit the learned text-3D correspondence to help ground the text queries to the referred 3D objects without regarding to 2D images. Through the designed scheme, a significant breakthrough is achieved than previous works, and the advantage of our 3D-VLA are also analyzed in detail. We believe these analyses can provide valuable insights to facilitate the future research of weakly supervised 3D visual grounding.